# Origin of effective graviton mass from spontaneous symmetry breaking of scalar field

### v1

@@ III Effective graviton mass through local U(1) SSB
Substituting this in action (6) leads to (details in Appendices A and B):
@@
This is the second key result of this work, and we want to discuss the following points. First, like in the previous case, the above action corresponds to extended FP action corresponding to the mass deformation parameter $\alpha=1/2$. Like in the previous case, the trace of the metric tensor is dynamical, and the metric tensor has six DoF with no ghost (details in Ref. [88]).
Second, by rewriting graviton mass in terms of the mass of the gauge field ($m_{A}$), we have:
@@
A curious reader might wonder: Can the aforementioned mass value be generated in a more restrictive model with fewer parameters, such as the Coleman-Weinberg model [81, 82], where $m_{\Phi}=0$? In the Coleman-Weinberg mechanism, SSB occurs due to radiative corrections. By repeating the above steps, we obtain the graviton mass as (details in Ref. [88]):
@@ IV Discussion
Our work rests on two robust frameworks: GR and quantum field theory. Our analysis demonstrates that non-perturbative phenomena, like SSB, indeed leave a significant imprint on gravity, and the effective graviton mass emerging as a natural consequence. Although our analysis is classical, the results remain valid since the energy scale in the SM of particle physics is small compared to the Planck energy scale. Therefore, we do not expect quantum corrections to alter our conclusions. While the analysis here primarily focuses on the complex scalar field, the analysis can be extended to any SSB mechanism.
@@
Given that all matter fields carry energy and are coupled to gravity, it is natural to ask: Do non-perturbative phenomena, like SSB, impact gravity? What effect do masses generated via SSB have on gravity? Typically, gravity is considered irrelevant at these energy scales and is neglected in such phenomena [97, 98]. While this assumption holds true perturbatively, our findings show that its validity becomes uncertain when non-perturbative effects like SSB are considered. We have demonstrated that SSB can produce non-trivial infrared (IR) effects on gravity. Specifically, we showed that gravitons acquire an effective mass due to the SSB of the global or local symmetry of a complex scalar field minimally coupled to gravity.
In this context, it is essential to emphasize that SSB involves concealing symmetry through a gauge choice while the underlying gauge symmetry remains unbroken, as stipulated by Elitzur’s theorem [99]. In gauge theory, physical observables are gauge-invariant and are typically computed through functional integrals over all gauge field configurations. However, to avoid overcounting due to the equivalence of any two points on a gauge orbit, a gauge-fixing procedure that intersects each gauge orbit precisely once is necessary. The manifestation of SSB, resulting from the non-trivial vacuum expectation value of the order parameter in the matter sector, represents a form of gauge choice. While this confers mass to gauge bosons, the intrinsic gauge symmetry persists through the Stueckelberg mechanism (see Secs. II and IV in [88]). Consequently, the effective mass gained by gravitons through the SSB mechanism does not imply breaking diffeomorphism invariance.
@@
In this work, we have explicitly demonstrated that the spontaneous symmetry breaking of local and global symmetry in the matter Lagrangian density leads to a non-zero mass for gravitons. Furthermore, we have established that the massive graviton action corresponds to the extended FP action ($\alpha=1/2$). Interesting $\alpha=1/2$ leads to 6 DoF and no ghost. Among these, two are helicity-2 DoF, two are helicity-1 DoF, and two are scalar DoF. Interestingly, the scalar DoF is coupled to the trace of the energy-momentum tensor. For radiation, the trace of the energy-momentum tensor is zero; hence, the coupling will be absent. Consequently, the extra scalar degree of freedom can effectively delineate these two distinct eras through the coupling. This represents a non-trivial aspect of the massive gravity theory that is not present in the theory of massless helicity-2 fields. Such a phenomenological model was studied to explain the acceleration of the current Universe [100, 101]. It will be interesting to investigate the cosmological consequence of such an analysis.
@@
Appendix A Effective graviton mass about Minkowski space-time from $U(1)$ symmetry breaking

### v2

@@ IV Effective massive spin-2 fields through local U(1) SSB
Substituting this in action (6) leads to:
@@
This is the second key result of this work, and we want to discuss the following points. First, like in the previous case, the above action corresponds to extended FP action corresponding to the mass deformation parameter $\alpha=1/2$. Like in the previous case, the trace of the metric tensor is dynamical, and the metric tensor has six DoF with no ghost (details in Appendices C and D).
Second, by rewriting the mass of spin-2 fields in terms of the mass of the gauge field ($m_{A}$), we have:
@@
Lastly, to understand the contribution of the 4-rank tensor $\mathscr{O}_{\mu\nu}^{\alpha\beta}$ w.r.t the Ricci scalar and tensor during the above SSB scenario. In the case of $\Lambda$CDM model the background Einstein’s equation leads to $R\sim\Lambda$. Hence, Ricci scalar is $\sim 10^{-64}\text{eV}^{2}$. On the other hand, the SSB in the dark matter sector (in our case) leads to mass squared of spin-2 field $m^{2}$ to be $10^{-56}\text{eV}^{2}$ which is $10^{8}$ order higher in magnitude. Substituting Eq. (18) in Eq. (6), we see that effective graviton mass is $10^{-56}\text{eV}^{2}$; hence, the contribution is from the fourth-rank tensor. Note that the concept of mass for the spin-2 field is well-defined only if the background metric solution is asymptotically flat since the action for the spin-2 field asymptotically approaches the extended Fierz-Pauli action ($\alpha=\frac{1}{2}$) at spatial infinity [111].
A curious reader might wonder: Can the aforementioned mass value be generated in a more restrictive model with fewer parameters, such as the Coleman-Weinberg (CW) model [87, 88], where $m_{\Phi}=0$? In the CW mechanism, the model exhibits one fewer parameter and is classically scale invariant. Furthermore, the Higgs mass is directly predicted. In the rest of this section, we adopt this approach and investigate the possibility of generating the effective mass of the spin-2 field. Including the radiative quantum corrections, the one-loop effective potential is:
$V_{eff}(\phi_{c})=-\frac{3g_{A}^{4}}{32\pi^{2}}\phi_{c}^{4}\Big{[}1-2\log\left(\frac{\phi_{c}^{2}}{v^{2}}\right)\Big{]}\,,$ (21)
The above potential has a maximum at $\phi_{c}=0$, whereas a minimum at $\phi_{c}=v$. Thus, the effective action of scalar quantum electrodynamics at one-loop level can be expressed as
$S_{QED}^{(1-lopp)}=\int d^{4}x\sqrt{-g}\,\Big{[}-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-(\mathscr{D}_{\mu}\phi)^{\dagger}(\mathscr{D}^{\mu}\phi)-V_{eff}(\phi)\Big{]}.$ (22)
Considering the energy-momentum tensor and the 4-rank tensor $\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma}$ associated with the above action, we obtain the following effective action of spin-2 field at the quadratic level
$\begin{split}S_{h}&=\int d^{4}x\sqrt{-\bar{g}}\Big{[}-\frac{1}{2}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{2}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h+\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}\\ &-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}+\kappa^{2}V_{eff}(\phi_{c}=v)\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\Big{]},\end{split}$ (23)
where
$V_{eff}(\phi_{c}=v)=-\frac{3g_{A}^{4}}{32\pi^{2}}v^{4}=-\frac{3}{32\pi^{2}}m_{Z}^{4}=-\frac{M^{2}v^{2}}{8}.$ (24)
By repeating the above steps, we obtain the graviton mass as:
@@ V Discussion
Our work is based on two well-established frameworks: GR and quantum field theory. Our analysis demonstrates that non-perturbative phenomena, such as spontaneous symmetry breaking (SSB), can significantly impact gravity, leading to an effective mass for spin-2 fields as a natural outcome. Although our analysis is classical, the conclusions remain robust because the energy scale of the Standard Model (SM) in particle physics is much lower than the Planck scale. As a result, quantum corrections are unlikely to alter our conclusions. While our study primarily focuses on a complex scalar field, the approach can be extended to any SSB mechanism.
@@
Given that all matter fields carry energy and interact with gravity, it is natural to ask: Do non-perturbative phenomena like SSB affect gravity? How do masses generated through SSB influence gravity? Traditionally, gravity is considered negligible at these energy scales and is often ignored in such contexts [112, 113]. While this assumption holds for perturbative effects, our results indicate that it may not apply when non-perturbative effects like SSB are considered. We have shown that SSB can induce non-trivial infrared (IR) effects on gravity. In particular, we demonstrated that spin-2 fields can acquire an effective mass through the SSB of a global or local symmetry in a complex scalar field minimally coupled to gravity.
As mentioned earlier, it is crucial to highlight that Einstein’s field equations are highly non-linear, allowing solutions only in a few highly symmetric cases. The perturbative approach discussed here is valuable because solving the Einstein equations in full generality is often not feasible. However, before performing a perturbative expansion, it is necessary to establish a background—a manifold that represents either the average universe or the background space-time. In General Relativity, this averaging process is challenging because it involves dealing with complex non-linear interactions. The impact of these non-linearities on the overall expansion is known as backreaction [89, 90]. Our approach, which does not rely on any assumptions about the background metric, is particularly effective due to the non-linear nature of the potential. Even without precise knowledge of the background metric, our method effectively captures the influence of the non-linear potential on perturbations, especially in the context of SSB. As mentioned earlier, since the background metric $\bar{g}_{\mu\nu}$ is chosen to be arbitrary, it has six independent components, which also implies that the resultant metric perturbation $h_{\mu\nu}$ has 6 DoF which is mathematically consistent with our analysis. On the other hand, it is also important to note that in asymptotic spatial infinity, the DoF of metric perturbation reduces to 2 as the energy-momentum vanishes there identically. As a result, there would not be any SSB in the matter sector.
In Refs. [114, 115], it is demonstrated that massive gravity permits a well-defined accelerating solution exclusively within open FLRW universes, and even in such contexts, perturbation theory reveals instabilities. In contrast to the aforementioned studies, which treat massive gravity as a modification of gravitational theory in cosmology, this work begins with GR and demonstrates that the non-zero mass of metric perturbations emerges only in the presence of SSB within the matter sector of the theory, within the framework of linearized gravity. Given that the initial gravitational theory is GR, the linearized massive gravity theory derived following the consideration of SSB explains the late-stage accelerating phase of the Universe, assuming it to be spatially flat due to the degravitation effect, as discussed in detail previously. Furthermore, it is important to emphasize that, in deriving the linearized massive gravity theory, we did not adopt dRGT massive gravity as the foundational theory, as is the case in Refs. [114, 115].
It is important to recognize that gauge transformations are not physical symmetries, as they do not correspond to any intrinsic property of a physical system or represent physical charges. Instead, they stem from the redundancy in the formalism. On the other hand, the Higgs mechanism is a well-known method for spontaneously breaking the physical symmetry of a system. However, in the Higgs mechanism, it is not the gauge symmetry itself that is broken but rather the symmetry of the potential. In this context, it is crucial to highlight that SSB involves concealing the symmetry through a gauge choice while the underlying gauge symmetry remains intact, as stated by Elitzur’s theorem [116, 117, 118, 119]. In gauge theory, physical observables are gauge-invariant and are usually computed via functional integrals over all gauge field configurations. To avoid overcounting equivalent points on a gauge orbit, a gauge-fixing procedure is employed to ensure that each gauge orbit is intersected precisely once.
The manifestation of SSB, characterized by the non-trivial vacuum expectation value of the order parameter in the matter sector, effectively represents a gauge choice. While this process grants mass to gauge bosons, the underlying gauge symmetry remains preserved via the Stueckelberg mechanism (see Appendix C). Consequently, the effective mass acquired by spin-2 fields through the SSB mechanism does not imply the breaking of diffeomorphism invariance.
For instance, in the case of local $U(1)$ symmetry breaking in scalar quantum electrodynamics, the non-trivial vacuum expectation value arising from SSB in the field potential leads to a non-zero mass for the $U(1)$ gauge field. This mass is physical and can be observed through the Meissner effect in superconductors (see [120]). While it may seem that the gauge symmetry is broken since the gauge field acquires a mass term at the action level after SSB, this is not the case, as Elitzur’s theorem ensures that the underlying gauge symmetry remains intact, as discussed earlier. In our scenario, diffeomorphism invariance is similarly preserved for any arbitrary background.
@@
In this work, we have explicitly demonstrated that the spontaneous symmetry breaking of local and global symmetry in the matter Lagrangian density leads to a non-zero mass for spin-2 fields. Furthermore, we have established that the massive spin-2 field action corresponds to the extended FP action ($\alpha=1/2$). Interesting $\alpha=1/2$ leads to 6 DoF and no ghost. Among these, two are helicity-2 DoF, two are helicity-1 DoF, and two are scalar DoF. Interestingly, the scalar DoF is coupled to the trace of the energy-momentum tensor. For radiation, the trace of the energy-momentum tensor is zero; hence, the coupling will be absent. Consequently, the extra scalar degree of freedom can effectively delineate these two distinct eras through the coupling. This represents a non-trivial aspect of the massive gravity theory that is not present in the theory of massless helicity-2 fields. Such a phenomenological model was studied to explain the acceleration of the current Universe [121, 122] in terms of degravitation (see Appendix D). It will be interesting to investigate the cosmological consequence of such an analysis.
@@
Appendix A Effective mass generation of spin-2 field in arbitrary curved spacetime